4. Evaluate (2 + y)dA, where D is the triangle with vertices (0,0), (0,1),(1,0).
Answer #1

I= 11 (**)dla 1 (221G2) (0,1) Jo y - y = Y2Y (1-21) 249-21 (0,0) 7-0 = 1-0 (3-1) one couro Jo xar-y (110) 9 -0 = 1-0 (2-0) Co